OpenStudy (anonymous):

Which of the following represents the zeros of f(x) = x3 − 14x2 + 61x − 84?

OpenStudy (anonymous):

@lyn

OpenStudy (anonymous):

@LynFran

OpenStudy (anonymous):

@LynFran

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

try some of the factors of 84 f(4) = 0 so (x -4) is a factor

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

3 a factor?

OpenStudy (welshfella):

now as -84 is the last term and x - 4 is a factor you could see if the factors of 21 could be zeros try finding f(3) and f(7)

OpenStudy (anonymous):

so7,4, and 3

OpenStudy (welshfella):

yes 3 is a factor

OpenStudy (welshfella):

3 is a zero i meant

OpenStudy (anonymous):

yes, may u help on another?

OpenStudy (welshfella):

well is f(7) = 0 ?

OpenStudy (welshfella):

you need to check that f(3) and f(7) = 0

OpenStudy (anonymous):

ok, is 3 a factor?

OpenStudy (anonymous):

or -3

OpenStudy (welshfella):

i checked f(3) it is = 0 so 3 is a zero

OpenStudy (anonymous):

and i checked 7 and it is a zero

OpenStudy (welshfella):

3 is not a factor (x -3) will be a factor

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

may you help on a nother

OpenStudy (welshfella):

f(7) also = 0 yes 4,3 and 7 are correct

OpenStudy (anonymous):

Which of the following is a polynomial with roots 3, 5i, and −5i?

OpenStudy (welshfella):

ok

OpenStudy (welshfella):

if x = 3 then one factor will be (x - 3) and the other 2 will be (x + 5i) and (x - 5i) what do you get when you multiply x + 5i and x - 5i ?

OpenStudy (anonymous):

x^2-25i

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

not quite (x - 5i)(x + 5i) = x^2 + 5xi - 5xi -5i*5i = x^2 - 25i^2 = x^2 - 25* -1 = x^2 + 25 remember that i^2 = -1

OpenStudy (anonymous):

ahh ok then what?

OpenStudy (welshfella):

so to get your polynomial you must multiply:- ( x - 3)(x^2 + 25)

OpenStudy (anonymous):

then we multiply by (X-3) for the answer right?

OpenStudy (anonymous):

yeah that right

OpenStudy (anonymous):

@welshfella what is our final answer?

OpenStudy (welshfella):

well do the multiplication (x - 3)(x^2 + 25) = x^3 + 25x can you continue?

OpenStudy (welshfella):

you need to multiply the second parenthesis by -3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-3x^2-75

OpenStudy (anonymous):

f(x) = x^3 − 3x^2 + 15x − 75 IS THAT THE ANSWER?

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

its 25x not 15x

OpenStudy (anonymous):

sry i mistyped

OpenStudy (anonymous):

thanks dude

OpenStudy (welshfella):

yw